Hi! My question is: What are the dimensions of a rectangle having perimeter 28 cm and a diagonal of length 10 cm. I'm hoping that using the pythagorean theorem is the correct way to go about this, but I keep getting strange answers. Any help would be greatly appreciated! Thanks!
Answer by richard1234(7193)

Using the Pythagorean theorem works well. If you can intuit the problem, recognize the 6-8-10 right triangle and note that 2(6+8) = 28, that works even better.

If you were to use the Pythagorean theorem, then you could say that

given the constraint that x+y = 14. Then, you could say that y = 14-x, replace to obtain the equation



Then solve for x and y.